40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  QIL elog  Not logged in ELOG logo
Message ID: 1199     Entry time: Thu Dec 9 02:53:22 2010
Author: Koji 
Type: Laser 
Category: GYRO 
Subject: Double Hacky? Trio? Even Hacky quartet! 

[Koji, Zach, Alastair, Frank]

After our very hacky trials, we found the noise source in the low frequency band:
It is the air current on the input/output optics even though the mechanism is still unknown.

We need:

- to know why the air flow in the I/O optics causes the noise.
- to make firm wind shield for the input/output optics too.


First Zach, Alastair and Koji started the afternoon experiment.

1. The first attempt was to put the Al foil on the table in the box to prevent the air flow from the table holes.
We practically saw no improvement.
Frank later pointed that the holes of the table are individually sealed. Therefore it was reasonable not to see any improvement.

2. The second attempt was to introduce He gas into the box just in order to see any possible change in the noise spectrum.
We went to the biology stock room to get the tubing for the gas introduction. As soon as we introduced the gas into the box,
the optical path length started to show drastic change (doe to reflactive index of He?). At the same time we started to see vertical
misalignment. We no longer could lock the cavity on TEM00. It may be caused by the temperature change in the box as the
introduced He was indeed quite cold due to adiabatic expansion? We also observed our voices got strange as we expected.

The He cyrinder is still in the lab. We may prepare some reservoir to accumulate the gas and let it heat???

At this point Zach and Alastair left the lab.

3. The third attempt was to put the air baffles made by Al foils along the cavity path. It made no change in the noise.

4. I was checking the PDH box. I found that the phase setting of the VCO loop was at the edge (10.00). I added a short BNC in the PD path.
Now the phase for the maximum optical gain is 9.60. The phase was adjusted such that the servo starts oscillation with as lower gain as possible.

I was observing the VCO feedback (the current gyro signal). It has ~100mVpp 40MHz signal coupled. It turned out that this signal comes from the
DAQ. I put a 5MHz LPF between the PDH box out mon and the DAQ cable.

The signal is also sent to SR560 to filter out the high freq fluctuations which makes the observation quite tricky.

5. I found that if the PZT output goes to negative on the oscilloscope, the gyro signal have larger noise than usual.
I put an offset to the slow servo such that it does not go into the negative side.

6. As the noise below 100Hz has non-stationary behavior I tried to find any sign of scatteriing / clipping. No luck.

Then I started the attept of putting Al foils to cover the input/output optics.
It indeed helped to reduce the noise between 0.1-20Hz! The improvement is factor of ~5 between 1~10Hz.
See the plot and the first photo. This was one of the hackiest work I ever had...

7. As I could not believe this improvement by myself, I told the story to Frank who was in the room at this time.
He got excited and proposed several ideas.

7a. First we removed the Al foils in order to confirm the effect is real. Yes it is real.

7b. The imptovement could have been given by settling the air flow or shadowing the ambient light.
Frank pointed out that our PD is sensitive to the visible wavelength. He brought the VIS light blocking
filters from PSL lab. The filters didn't help the noise level.
The noise from the ambient light was rejected at least for now.

7c. We partially put the Al foils on the table. The effect was not quite localized but relatively strong at the downstream side of the input optics.
There is the air conditioning working above those optics. We put the Al foils to block the A.C. air. It improved the noise level by factor of 2-3.
Anyway, we confirmed that the main cause is the air flow.

7d. Frank found the plexiglas plates at the edge of the room. We constructed the temporary wind shield by them. See the second photo.
It improved the noise level as shown in the plot, although the effect was not as good as the Al foil case.

8. We decided to finish the work of the day. We cleaned up our mess (expect for the foil on the table in the box).
The stray Al foils were disposed. The Al reel was returned to the suspension lab.

 

Attachment 1: Gyro_Spectra_101208.pdf  133 kB  Uploaded Thu Dec 9 00:09:30 2010  | Show | Hide all | Show all
Attachment 2: 100_1158.jpg  1.367 MB  Uploaded Thu Dec 9 00:10:16 2010  | Hide | Hide all | Show all
100_1158.jpg
Attachment 3: 100_1162.jpg  1.199 MB  Uploaded Thu Dec 9 00:11:00 2010  | Hide | Hide all | Show all
100_1162.jpg
ELOG V3.1.3-